hi i was wondering what the numerical pattern is for 7,19,31,43
Found 2 solutions by oberobic, jim_thompson5910:
Answer by oberobic(2304)

43-31 = 12
31-19 = 12
19-7 = 12
.
So it appears the sequence is adding 12 to whatever number you started with.

Answer by jim_thompson5910(35256)

You're adding 12 each time.

How do I know that? Well after a while you get used to seeing patterns like this, but notice how 19-7 = 12, 31-19 = 12, and 43-31 = 12

If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim